# PyTomography: A Python Library for Medical Image Reconstruction Luke Polson1,2, Roberto Fedrigo1,2, Chenguang Li1,2, Maziar Sabouri1,2, Obed Dzikunu2,3, Shadab Ahamed1,2, Nicolas Karakatsanis4, Arman Rahmim1,2,3,5,∗, Carlos Uribe2,5,6,∗

### v3

@@ VI.B. Validation of PyTomography
    Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617. Simulated data was reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography. Coronal slices and 1D profiles of reconstructions are shown in Figure 7 and are nearly identical. The image noise, quantified by the standard deviation of predicted counts in the liver, was 0.789 counts (STIR) and 0.793 counts (PyTomography). The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography). The reconstruction time in PyTomography (5.4 s) was significantly faster than the time required in STIR (51542.2 s), owing to the speed-up factor of GPU implementation.
@@
    Clinical SPECT Validation With Vendor: A Jaszczak phantom was filled with Lu-177 and scanned at 180 projections ($128\times 128$ with a resolution of $4.8\text{mm}\times 4.8\text{mm}$) using a Siemens Symbia T Series SPECT/CT scanner with Medium Energy collimators. Reconstruction was performed with OSEM (4 iterations, 10 subsets) using (i) the manufacturer scanner software and (ii) PyTomography. Reconstructions and corresponding 1 dimensional profiles are shown in Figure 8; they appear similar. 3D slicer was used to obtain sample segmentations of the hot and cold regions; the mean source-to-background ratios were 20.97 (vendor) and 21.76 (PyTomography). The image noise, quantified by the standard deviation of predicted counts in the cold region, were 0.217 counts (vendor) and 0.178 counts (PyTomography).
@@
    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image34 used in the PET example (556 million events) was reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR. Resulting reconstructions are shown in Figure 9. The mean activity ratio in grey to white matter was 2.05 (PyTomography) and 2.03 (CASToR). The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR). Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s per iteration for reconstruction. CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores. The reconstruction was performed with 24 CPU cores and reconstruction required approximately 455 s per full iteration.
@@
Figure 8: Sample axial slices (top) and 1D profiles (bottom) of a 177Lu Jaszczak phantom reconstructed using OSEM (4it/10ss) with attenuation/PSF modeling and triple energy scatter correction. Reconstructions were performed using (i) the scanner software, and (ii) PyTomography. 1D profiles correspond to the horizontal lines shown on the axial slices.

### v4

@@ Appendix A Validation of PyTomography
    Monte Carlo SPECT Validation With STIR: Radioactivity concentrations in a digital phantom were selected to correspond to a 24-hour post injection prostate cancer patient (1700 MBq total activity) imaged via the prostate-specific membrane antigen (PSMA)-targeting radiopharmaceutical, 177Lu-PSMA-617. SPECT acquisition was simulated in SIMIND with dimensions $128\times 384$, a pixel size of $0.48~{}\text{cm}\times 0.48~{}\text{cm}$, 120 projections, an acquisition time of 15 s/projection, and use of Siemens medium energy collimators. Data were reconstructed with OSEM (2 iterations, 8 subsets) using (i) STIR and (ii) PyTomography. Coronal slices and 1D profiles of reconstructions are shown in Figure 5 and are nearly identical. The image noise, quantified by the standard deviation of predicted counts in the liver, was 0.789 counts (STIR) and 0.793 counts (PyTomography). The mean predicted counts in the kidneys were 1.764 (STIR) and 1.761 (PyTomography). The reconstruction time in PyTomography (5.4 s) was significantly faster than the time required in STIR (51542.2 s), owing to the speed-up factor of GPU implementation.
@@
    Clinical SPECT Validation With Siemens and MIM: A NEMA phantom was filled with 177Lu at a 10:1 source to background ratio and scanned at 96 projections ($128\times 128$ with a resolution of $4.8\text{mm}\times 4.8\text{mm}$) for 15 s per projection using a Siemens Symbia T Series SPECT/CT scanner with Medium Energy collimators. Reconstruction was performed with OSEM (4 iterations, 8 subsets) with attenuation/PSF modeling and TEW scatter correction, using (i) the manufacturers software: Siemens Symbia Intevo Excel v9.0.31.3, (ii) MIM SPECTRA v7.2.1, and (iii) PyTomography. Reconstructions and corresponding 1 dimensional profiles are shown in Figure 6; the PyTomography and Siemens reconstructions are qualitatively more similar than the MIM reconstruction. 3D slicer was used to obtain sample segmentations of (i) the largest NEMA sphere, (ii) the central cold portion, and (iii) the warm background region. The mean counts in the largest NEMA sphere were 14.07 (Siemens), 12.96 (MIM), and 13.49 (PyTomography). The mean counts in the central cold portion were 0.732 (Siemens), 0.704 (MIM), and 0.688 (PyTomography). The image noise, quantified by the standard deviation of predicted counts in the warm region, were 0.196 (Siemens), 0.168 (MIM), and 0.179 (PyTomography).
@@
    Monte Carlo PET Validation With CASToR: The ultra-high resolution 18F-FDG PET/ T1-weighted MRI brain image [25] used in the PET example was simulated in GATE without random, scatter, attenuation, or positron range modeling to obtain a ultra high count dataset consisting only of primary events. The data were reconstructed using LM TOF OSEM (2 iterations, 21 subsets) with both PyTomography and CASToR. Resulting reconstructions are shown in Figure 7. The mean activity ratio in grey to white matter was 2.05 (PyTomography) and 2.03 (CASToR). The noise in the image was estimated as the standard deviation of counts in white matter divided by the mean activity in white matter; the values were 0.329 (PyTomography) and 0.321 (CASToR). Apart from processing the ROOT files and building the sensitivity image, the time required for reconstruction in PyTomography was 14 s to build the system matrix and required 59 s/it for reconstruction. CASToR reconstruction was ran on a separate HP Z8 Computer Desktop with 2 Intel Xeon Gold 6136 CPUs at 3.00 GHz with a total of 48 CPU cores. The reconstruction was performed with 24 CPU cores and reconstruction required approximately 455 s per full iteration.
@@
Figure 6: Sample axial slices (top) and 1D profiles (bottom) of a 177Lu NEMA phantom reconstructed using OSEM (4it/8ss) with attenuation/PSF modeling and triple energy scatter correction. Reconstructions were performed using (i) the Siemens scanner software, (ii) MIM SPECTRA Reconstruction, and (iii) PyTomography. 1D profiles correspond to the horizontal lines shown on the axial slices.